On September 10, a credit card account had a balance of $490. Purchases of $47.32 were made on September 12, $157.55 was charged on September 20, and $57.73 was charged on September 30. A payment of $175 was made on September 28. The annual interest on the average daily balance is 21.25%. Find the finance charge (in dollars) on the October 10 bill. (Use 365 for the number of days in a year. Round your answer to the nearest cent.)
Answer #1
Please see the below table for credit card account
Date Payment or Purchase Balance each day No. of days untill balance charge Unpaid balance times no. of days
September 10 to 11 490 2                           980.00
September 12 to 19 47.32 537.32 8                       4,298.56
September 20 to 27 157.55 694.87 8                       5,558.96
September 28 to 29 -175 519.87 2                       1,039.74
September 30 to October 9 57.73 577.6 10                       5,776.00
30                     17,653.26
Therefore,
Average Daily Balance = (sum of the total amount owned each day of month) / No. of days in the billing period
Average Daily Balance = 17653.26/30
Average Daily Balance = $588.44
The Finance charge is
I = Prt
I = 588.44*1.77%*1
I = $10.41
the finance charge on October 10, bill is $10.41
r (rate) is given 21.25% annually so we have to convert into monthly
monthly rate = 21.25/12 = 1.77
t = time 1 month